OpenStudy (domebotnos):

convert r=2sin(theta)-4cos(theta) to a rectangular equation

OpenStudy (bradely):

r=2sin(theta)-4cos(theta) cos theta =x/r, and sin(theta) = y/r r^2 =x^2+y^2 r= 2(x/r)-4(y/r) r^2 =2x-4y x^2+y^2 =2x-4y x^2-y^2-2x+4y=0 Source: http://www.mathskey.com/question2answer/ask

OpenStudy (bradely):

Please see example problem here. http://www.mathskey.com/question2answer/16710/identify-the-graph-of-r-4-cos-theta-4-sin-theta

OpenStudy (ikram002p):

r^2=x^2+y^2 x=r cosθ → x/r=cosθ→ x/\(\sqrt(x2+y\)2)=cosθ y=r sinθ → y/r=sinθ→ y/\(\sqrt (x2+y2)\) =sinθ